OpenStudy (anonymous):

This one is not that obvious to differentiate. Help please? f(u)=lnu / 1 + ln2u

OpenStudy (anonymous):

quotient rule for this one

OpenStudy (anonymous):

\[\left(fg\right)'=\frac{gf'-fg'}{g^2}\] with \[f(x)=\ln(x), f'(x)=\frac{1}{x},g(x)=1+\ln(2x),g,(x)=\frac{1}{x}\]

OpenStudy (anonymous):

Im an idiot.... Over thinking as usual...Thank you satellite73!!

OpenStudy (anonymous):

yw lots of ugly algebra i imagine